# From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy

### v1

@@ Abstract
Problem-solving therapy (PST) is a structured psychological approach that helps individuals manage stress and resolve personal issues by guiding them through problem identification, solution brainstorming, decision-making, and outcome evaluation. As mental health care increasingly integrates technologies like chatbots and large language models (LLMs), understanding how PST can be effectively automated is important. This study leverages anonymized therapy transcripts to analyze and classify therapeutic interventions using various LLMs and transformer-based models. Our results show that GPT-4o achieved the highest accuracy (0.76) in identifying PST strategies, outperforming other models. Additionally, we introduced a new dimension of communication strategies that enhances the current PST framework, offering deeper insights into therapist-client interactions. This research demonstrates the potential of LLMs to automate complex therapeutic dialogue analysis, providing a scalable, efficient tool for mental health interventions. Our annotation framework can enhance the accessibility, effectiveness, and personalization of PST, supporting therapists in real-time with more precise, targeted interventions.
From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy
@@ 1 Introduction
In recent years, the increasing complexities of modern life have significantly influenced the prevalence of developing mental health conditions. According to the American Psychological Association (APA), the COVID-19 pandemic, global conflicts, racism, inflation, and climate disasters have led to widespread collective trauma among Americans, with increased chronic illnesses in those aged 35 to 44 and a significant rise in mental health diagnoses (APA, 2023). In response, the World Health Organization (WHO) underscores the urgent transformation of global mental health care (WHO, 2022). Building on this, Problem-Solving Therapy (PST) has emerged as a pivotal therapeutic approach to help individuals manage these stressors and challenges more effectively (Song and Sun, 2019). Rooted in the principle that improving problem-solving skills reduces psychological distress, PST fosters resilience and well-being (Arean et al., 2008). By guiding individuals to address immediate problems with practical strategies (Figure 1), PST strengthens coping mechanisms and mitigates negative impacts on both mental and physical health (Gellis and Kenaley, 2008).
Figure 1: A real-life example of therapist and client exchanges in a problem solving therapy session.
As rapidly evolving trends continue to shape the critical high-stakes area of mental health, there is currently a lack of systematic methods for exploring and understanding PST between clients and therapists Chiu et al. (2024). This gap becomes particularly relevant with the increasing prevalence of mental health chatbots, especially those deployed using large language models (LLMs) Lawrence et al. (2024); Cho et al. (2023). Before these models are widely adopted, it is crucial to examine how mental health as well as PST strategies are implemented in practice. Previous research highlighted the variability in therapists’ application of PST techniques, which are often tailored to individual client needs and contexts Andrade-Arenas et al. (2024). In PST sessions, therapists follow a structured process: identifying the problem, brainstorming solutions, making decisions, and evaluating outcomes—all within a supportive framework. This approach equips clients with PS skills to tackle future challenges, enhancing their sense of control and overall coping abilities (Malouff et al., 2007). Therefore, while the new technological solutions hold promises, it is essential to understand how PST can be effectively automated, ensuring these tools are responsive and sensitive to the varied situations of clients Izadi and Forouzanfar (2024).
@@
Our work addresses this gap by using LLMs to automate the annotation of PST, leveraging both closed and open-weight models to categorize therapist utterances. This approach allows us to efficiently analyze therapist-client interactions and enhances the scalability and precision of mental health interventions through automated systems. More specifically, we used anonymized real-life 240 PST conversations between family caregivers and their therapists, consisting of 68,306 dialogue exchanges. We applied both open- and closed-weight LLMs to automatically annotate therapist utterances with strategies based on existing and adapted codebooks of PST and communication strategies. Our results show that using LLMs for analyzing PST-related texts is efficient, as they achieve high accuracy for annotating PST strategies in therapeutic conversations. Additionally, our analysis is validated by examining the specific strategies therapists employed across multiple client sessions, highlighting how these strategies evolved as therapy progressed. Our approach not only facilitates a deeper understanding of therapist-client interactions but also supports the development of more responsive and adaptive mental health interventions by enabling the real-time identification and application of effective therapeutic techniques.
Our work contributes to the field of NLP and mental health care by introducing a framework for automating complex therapeutic dialogue analysis, providing a scalable, efficient tool for mental health interventions. In addition, using anonymized transcripts from real-world therapy sessions, our results advance PST research and demonstrate the transformative potential of LLMs in mental health care. Finally, adding a new dimension of communication strategies enables a deeper understanding of therapist-client interactions, thereby refining the existing PST framework and enhancing the potential for more targeted therapeutic interventions.
@@ 2.1 LLMs for Mental Health Care
LLMs have been applied for mental health care, leveraging distinct features such as conversational capabilities Ma et al. (2023); Zheng et al. (2023), content generation Yang et al. (2024); Smith et al. (2023), and detection of information patterns Posner et al. (2011); Mauriello et al. (2021). Based on 34 papers out of 313 publications from six databases Hua et al. (2024) identified three key application areas of LLMs in mental health care: 1) conversational agents for digital interventions, 2) resource enrichment for training and technical development, and 3) diagnosis of mental disorders and their sources. LLMs have been integrated into conversational agents (Bendig et al., 2022) and used for personal digital companions (Ma et al., 2023), emotional support (Zheng et al., 2023), and online counseling (Sharma et al., 2023; Filienko et al., 2024).Researchers applied LLMs to generate explanations for mental health analysis Yang et al. (2024), help people learn mindfulness Kumar et al. (2023), and create hypothetical case vignettes to social psychiatry (Smith et al., 2023). LLMs have also been used in classification of subtypes of suicide (Posner et al., 2011), and identification of sources of stress span (Mauriello et al., 2021).
However, gaps remain in previous research in prioritizing privacy, safety, and ethics (Hua et al., 2024; Wang et al., 2021a; Goel et al., 2023). For example, among 36 datasets identified in Hua et al. (2024), 55.6% were collected from social media, 13.9% were simulated by clinicians, crowd workers, and LLMs, and only 8.3% came from real-life counseling conversations. Moreover, labels in existing datasets are typically not validated by experts (Hua et al., 2024; Harrigian et al., 2020). The absence of relevant evaluation frameworks brings further challenges. Every cognitive and interpersonal therapy technique has unique, specialized structure and requirement for evaluating its fidelity of delivery(Sharma et al., 2023; Mehta et al., 2022). Examining how mental health professionals deliver interventions through therapy techniques in different contexts is essential for addressing these gaps and improving professionals’ practices.
@@ 2.2 LLMs as (Co-)Annotators
LLMs have been increasingly used for annotation since acquiring high-quality labeled data is time-consuming and costly when relying solely on human experts for annotation Wang et al. (2021b, 2024); Bouzoubaa et al. (2024b). Methods such as zero-shot, without any in-domain training Liu et al. (2024); Reynolds and McDonell (2021); Bouzoubaa et al. (2024a) and few-shot learning, incorporating examples of the task Brown et al. (2020) have been used to optimize prompts Reynolds and McDonell (2021); Goel et al. (2023). Additional research showed LLMs outperformed crowd workers on annotation tasks including relevance, stance, topics, frame detection, and categorizing sentence segments Gilardi et al. (2023); He et al. (2023); Törnberg (2023). Wang et al. (2021b) found that GPT-3 cost 50% to 96% less than human for annotation. On the other hand, Zhu et al. (2023) and Ziems et al. (2024) showed that the performance of LLMs can vary substantially depending on tasks, datasets, and labels, claiming the irreplaceableness of human annotators. Wang et al. (2021b) and Li (2024) found that human-LLM collaborative approaches—either through human annotation of LLM-labeled data or combining both annotations—outperformed LLM-only or human-only approaches. Li et al. (2023c) further explored the allocation between humans and LLMs for annotation and contributed to a paradigm using uncertainty to estimate LLMs’ annotation capability. Researchers also developed human-LLM collaborative annotation systems to support the annotations Kim et al. (2024); Tang et al. (2024).
Researchers in healthcare explored a human-LLM collaboration using prompt engineering to extract medication information and found that LLMs accelerated this process with accuracy comparable to trained medical NLP annotators Goel et al. (2023). They also demonstrated the potential of a fine-tuned T5-based model for scaling data collection by generating annotation scores from interview transcripts for Bipolar Disorder and Schizophrenia Aich et al. (2024).
The studies so far, however, show limited progress on how LLMs perform as annotators in PST-related therapy. Investigating their performance on real-world, closed-domain data, like medical records, could offer valuable insights into their annotation capabilities. This experimentation may also highlight their potential beyond conversational agents, improving clinical workflows in mental health therapy delivery.
@@ 3.1 Data
Problem-Solving Therapy Codebook. The annotation codebook was developed based on the core strategies derived from the ADAPT model of PST, which includes five steps: 1) developing a positive Attitude, 2) Defining the problem and setting goals, 3) being creative and finding Alternative solutions, 4) Predicting how the solutions will work and making a plan, and 5) Trying out the solution plan Demiris et al. (2019). We refer to these as the ‘Problem-Solving (PS) core’ category. Upon further exploration of our data and various frameworks and codebooks for therapy conversations and strategies Li et al. (2023a); Ribeiro et al. (2013); Wang et al. (2023); Ghamari Kivi and Jamshiddoust Miyanroudi (2022), our domain experts suggested adding a new category of four strategies to this list, called ‘Communication’ strategies, to enhance the comprehensiveness of the analysis. These strategies are further explained in Table 6.
Data Annotation. To develop an evaluation set, we randomly selected 500 therapist utterances for detailed annotation. One domain expert alongside one researcher from our team performed the annotation. The annotators were instructed to select at most one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category, if applicable. If none of the strategies were suitable, we labeled it as ‘None’. We used Cohen’s Kappa to assess annotators’ agreement across different classes of PST strategies. Our findings show that the agreement scores ranged from 0.69 to 0.88, indicating a substantial agreement among the annotators, reflecting a reasonable level of consistency in identifying the strategies. The highest agreement was observed in classes that were more straightforward to identify, such as ‘Generating Alternative Solutions’, whereas the lowest agreement occurred in classes that were potentially more subjective or nuanced in nature, e.g., ‘Therapeutic Engagement’.
@@ 3.2 Classification of PST Strategies
LLM Prompting. LLMs, such as GPT-4, have been used for data annotation, offering capabilities for understanding and categorizing diverse datasets (Li et al., 2023a; Kuzman et al., 2023; Hoes et al., 2023). In this work, we employed these models to annotate therapist utterances, aiming to identify the PST strategy used in each case. Specifically, we used two versions of OpenAI’s models: gpt-4o-2024-08-06 (OpenAI, ) (referred to as GPT-4o) and gpt-4-0125-preview (OpenAI, 2024) (referred to as GPT-4), as well as two open-source models, Llama-3.1-70B-Instruct (Meta, ) (referred to as Llama) and Yi-1.5-34B (Young et al., 2024) (referred to as Yi). All models were deployed with a zero-temperature setting to ensure deterministic output, enhancing the consistency of our annotations (See Appendix C for hyperparameters). Using this approach, we compared the efficiency and accuracy of these models against our 500 annotated utterances. We used two approaches to test how well LLMs can annotate data:
    No Context: We provided specific instructions to LLMs, followed by definitions of each strategy and a few examples (one for each class). We then presented each therapist’s utterance for classification, without providing any additional conversational context.
@@
The labeling process with LLMs followed the same approach as with the annotators. In each prompt, we asked the models to select one strategy from the ‘PS core’ and ‘Communication’ categories or return ‘None’ if none applied. Prompts are available in Appendix D. We ran each LLM five times to test their stability for annotation. To evaluate the consistency of the annotation, we calculated the entropy of the labels generated by the LLMs (Li et al., 2023b). Entropy (Eq. (1)) measures the degree of uncertainty or disorder within a dataset. The higher the entropy value, the more uncertain the LLMs’ responses are.
@@
where $P_{\theta}(a_{ij}\mid p_{ij})$ is calculated as the frequency of a certain prediction among all predictions.
Model Training via LLM Labels. In addition to using LLM-based models, we also tested the performance of three transformer-based models to classify the utterances in our data. Using a subset of annotated data with the best performing LLM, we fine-tuned DeBERTa He et al. (2020), MentalBERT Ji et al. (2022), and FlanT5 Chung et al. (2024). Model details are provided in Appendix C. The choice of these models was driven by the need to explore a diverse range of transformer architectures while also including models tailored to mental health-related tasks. We fine-tuned each model for 3 epochs with a learning rate of 2e-5. Given that our annotation framework was based on two dimensions, ‘PS core’ and ‘Communication’ strategies, we fine-tuned two separate models for each dimension. Finally, we tested the performance of each fine-tuned model using our evaluation set of 500 human-annotated utterances.

### v2

# From Conversation to Automation: Leveraging LLMs for Problem-Solving Therapy Analysis

@@ Abstract
Problem-solving therapy (PST) is a structured psychological approach that helps individuals manage stress and resolve personal issues by guiding them through problem identification, solution brainstorming, decision-making, and outcome evaluation. As mental health care increasingly adopts technologies like chatbots and large language models (LLMs), it is important to thoroughly understand how each session of PST is conducted before attempting to automate it. We developed a comprehensive framework for PST annotation using established PST Core Strategies and a set of novel Facilitative Strategies to analyze a corpus of real-world therapy transcripts to determine which strategies are most prevalent. Using various LLMs and transformer-based models, we found that GPT-4o outperformed all models, achieving the highest accuracy (0.76) in identifying all strategies. To gain deeper insights, we examined how strategies are applied by analyzing Therapeutic Dynamics (autonomy, self-disclosure, and metaphor), and linguistic patterns within our labeled data. Our research highlights LLMs’ potential to automate therapy dialogue analysis, offering a scalable tool for mental health interventions. Our framework enhances PST by improving accessibility, effectiveness, and personalized support for therapists.
From Conversation to Automation: Leveraging LLMs for Problem-Solving Therapy Analysis
@@ 1 Introduction
In recent years, modern stressors like COVID-19, global conflicts, and climate disasters have led to rising mental health conditions, especially among adults 35-44 (APA, 2023). In response, the World Health Organization (WHO) calls for urgent transformation in mental health care (WHO, 2022). Problem-Solving Therapy (PST) helps manage these stressors by enhancing problem-solving skills, reducing psychological distress, and strengthening coping mechanisms (Song and Sun, 2019; Arean et al., 2008; Gellis and Kenaley, 2008).
Figure 1: A real-life example of therapist and client exchanges in a problem-solving therapy session.
As rapidly evolving trends continue to shape the critical area of mental health, systematic methods for analyzing PST interactions remain limited (Chiu et al., 2024). This gap is especially relevant with the rise of LLM-powered mental health chatbots (Lawrence et al., 2024; Cho et al., 2023). Before widespread adoption, it is crucial to assess how PST strategies are applied in practice, as therapist approaches vary based on client needs and contexts (Andrade-Arenas et al., 2024). PST follows a structured process—identifying problems, brainstorming solutions, decision-making, and evaluation—to strengthen clients’ problem-solving skills and coping abilities (Malouff et al., 2007). While new technologies hold promise, their adaptability and responsiveness must be ensured for effective PST automation (Izadi and Forouzanfar, 2024).
@@
Our work addresses this gap by using LLMs to automate PST annotation, using both closed- and open-weight models to categorize therapist utterances. We analyzed 240 real-world anonymized PST sessions (68,306 dialogue exchanges), applying LLMs to annotate strategies based on existing and adapted PST and Facilitative strategy codebooks. Our results show that LLMs efficiently and accurately annotate PST strategies. We also analyzed how these strategies evolve across sessions finding that therapists progressively shift from exploratory strategies, such as defining problems and brainstorming solutions, to strategies focused on implementation and evaluation. Additionally, we analyze therapeutic dynamics (autonomy, self-disclosure, metaphor) and linguistic properties in all utterances to understand how these strategies were applied. Our results show that therapists adapt their language to session goals—using more non-directive language and metaphors in early stages and transitioning to directive language and concrete communication as therapy progresses.
This approach enhances therapist-client interaction analysis, supports scalable and adaptive mental health interventions, and enables real-time identification of effective therapeutic techniques. Our framework advances NLP and mental health care, showing LLMs’ potential to transform PST research. Finally, integrating Facilitative strategies refines the PST framework, offering deeper insights into targeted therapeutic interventions.
@@ 2.1 LLMs for Mental Health Care
LLMs have been applied for mental health care, leveraging distinct features such as conversational capabilities Ma et al. (2023); Zheng et al. (2023), content generation Yang et al. (2024); Smith et al. (2023), and detection of information patterns Posner et al. (2011); Mauriello et al. (2021). Specifically, LLMs have been applied by conversational agents (Bendig et al., 2022) to serve as personal digital companions (Ma et al., 2023), provide emotional support (Zheng et al., 2023), and support online counseling (Sharma et al., 2023; Filienko et al., 2024). Researchers applied LLMs to generate explanations for mental health analysis Yang et al. (2024), help people learn mindfulness Kumar et al. (2023), and create hypothetical case vignettes to social psychiatry (Smith et al., 2023). LLMs have also been used in classification of subtypes of suicide (Posner et al., 2011), and identification of sources of stress span (Mauriello et al., 2021).
However, gaps remain in previous research in accessing data with high external validity to fully understand the potential LLMs for different mental health care settings (Hua et al., 2024; Wang et al., 2021a; Goel et al., 2023). For example, among 36 datasets identified in Hua et al. (2024), 55.6% were collected from social media, 13.9% were simulated by clinicians, crowd workers, and LLMs, and only 8.3% came from real-life counseling conversations. Moreover, labels in existing datasets are typically not validated by experts (Hua et al., 2024; Harrigian et al., 2020). The absence of relevant evaluation frameworks brings further challenges. Every cognitive and interpersonal therapy technique has a unique, specialized structure and requirement for evaluating fidelity of delivery(Sharma et al., 2023; Mehta et al., 2022). Examining how professionals deliver interventions through therapy techniques is essential for addressing these gaps.
@@ 2.2 LLMs as (Co-)Annotators
Acquiring high-quality labeled data is time-consuming and costly when relying solely on human experts for annotation Wang et al. (2021b, 2024); Bouzoubaa et al. (2024b). LLMs have been explored to support annotations through prompt engineering Reynolds and McDonell (2021); Goel et al. (2023); Bouzoubaa et al. (2024a), such as zero-shot, without any in-domain training Liu et al. (2024); Reynolds and McDonell (2021) and few-shot learning, incorporating examples of the task Brown et al. (2020). Research showed LLMs outperformed crowd workers on annotation tasks including relevance, stance, topics, frame detection, and categorizing sentence segments Gilardi et al. (2023); He et al. (2023); Törnberg (2023).
On the other hand, Zhu et al. (2023) and Ziems et al. (2024) showed that the performance of LLMs vary depending on tasks, datasets, and labels, claiming the irreplaceableness of human annotators. Wang et al. (2021b) and Li (2024) found that human-LLM collaboration—either through human annotation of LLM-labeled data or combining both annotations—outperformed LLM-only or human-only approaches. Li et al. (2023c) explored the allocation between humans and LLMs for annotation and contributed to a paradigm using uncertainty to estimate LLMs’ annotation capability. Researchers also developed human-LLM collaborative annotation systems to support the annotations Kim et al. (2024); Tang et al. (2024). Our study expands on previous research by employing LLMs to annotate therapy transcripts, thereby refining existing frameworks and models and informing broader applications of LLMs in mental health care research.
@@ 3.1 Data
Problem-Solving Therapy Codebook. In collaboration with PST experts, we developed the annotation codebook using the core strategies from the ADAPT model of PST, which includes five steps: 1) developing a positive Attitude, 2) Defining the problem and setting goals, 3) being creative and finding Alternative solutions, 4) Predicting how the solutions will work and making a plan, and 5) Trying out the solution plan Demiris et al. (2019). We refer to these as the ‘Problem-Solving (PS) Core’ category. Upon further exploration of our data and various frameworks and codebooks for therapy conversations and strategies Li et al. (2023a); Ribeiro et al. (2013); Wang et al. (2023); Ghamari Kivi and Jamshiddoust Miyanroudi (2022), our domain experts suggested adding a new category of four strategies to this list, called ‘Facilitative’ strategies, which we will use interchangeably with ‘Facilitators’ throughout the paper. These strategies are further explained in Table 6.
Data Annotation. To develop an evaluation set, we randomly selected 500 therapist utterances for detailed annotation. One domain expert alongside one researcher from our team performed the annotation. The annotators were instructed to select at most one strategy from the ‘PS Core’ category and one strategy from the ‘Facilitators’, if applicable. If none of the strategies were suitable, we labeled it as ‘None’. We used Cohen’s Kappa to assess annotators’ agreement across different classes of PST strategies. Our findings show that the agreement scores ranged from 0.69 to 0.88, indicating a substantial agreement among the annotators, reflecting a reasonable level of consistency in identifying the strategies. Agreement was higher for more straightforward classes (e.g., ‘Generating Alternative Solutions’) and lower for more subjective or nuanced classes (e.g., ‘Therapeutic Engagement’).
@@ 3.2 Classification of PST Strategies
LLM Prompting. LLMs, such as GPT-4, have been used for data annotation, offering capabilities for understanding and categorizing diverse datasets (Li et al., 2023a; Kuzman et al., 2023; Hoes et al., 2023). In this work, we employed these models to annotate therapist utterances, aiming to identify the PST strategy used in each case. Specifically, we used two versions of OpenAI’s models: gpt-4o-2024-08-06 (OpenAI, ) (referred to as GPT-4o) and gpt-4-0125-preview (OpenAI, 2024) (referred to as GPT-4), as well as two open-source models, Llama-3.1-70B-Instruct (Meta, ) (referred to as Llama) and Yi-1.5-34B (Young et al., 2024) (referred to as Yi-1.5). All models were deployed with a zero-temperature setting to ensure deterministic output, enhancing the consistency of our annotations (See Appendix C for hyperparameters). Using this approach, we compared the efficiency and accuracy of these models against our 500 annotated utterances. We used two approaches to test how well LLMs can annotate data:
    No Context: We provided specific instructions to LLMs, followed by definitions of each strategy and a few examples. We then presented each therapist’s utterance for classification, without providing any additional conversational context.
@@
Labeling with LLMs followed the same approach as with the annotators. In each prompt, we asked the models to select one strategy from the PS Core and Facilitative categories or return ‘None’ if applicable. Prompts are available in Appendix D. We ran each LLM five times to test their stability for annotation. To evaluate the consistency of the annotation, we calculated the entropy of the labels generated by the LLMs (Li et al., 2023b). Entropy (Eq. (1)) measures the degree of uncertainty or disorder within a dataset. The higher the entropy value, the more uncertain the LLMs’ responses are.
@@
where $P_{\theta}(a_{ij}\mid p_{ij})$ is calculated as the frequency of one prediction among all predictions.
Transformer-Based Models. In addition to using LLM-based models, we tested the performance of four transformer-based models—DeBERTa He et al. (2020), MentalBERT Ji et al. (2022), FlanT5 Chung et al. (2024), and ModernBERT Warner et al. (2024)—on classifying utterances. From the 14,417 utterances annotated by our best LLM, we randomly selected 5,000 utterances, excluding the evaluation set, and fine-tuned the models on this subset. The models were fine-tuned separately for the two annotation dimensions, PS Core and Facilitative categories, and evaluated on the 500-utterance human-annotated test set.
3.3 Therapeutic Dynamics
Previous section focused on identifying ‘which’ strategies are used. Here, we turn our attention to ‘how’ these strategies are applied by analyzing autonomy, self-disclosure, and metaphor usage.
Autonomy is a cornerstone of effective counseling, enhancing therapeutic alliances and treatment outcomes Dwyer et al. (2011). While its importance is well-established, measuring autonomy has largely focused on theory and behavior, overlooking linguistic markers. Previous research Shah et al. (2022) has often relied on open-ended versus close-ended questions as a proxy for assessing autonomy. However, we adopt a more comprehensive definition grounded in the literature to better capture the nuances of autonomy in therapeutic interactions. To address this, we define autonomy through three categories: directive, non-directive, Overholser (1987) and N/A. Directive approaches involve therapist-led guidance, useful but potentially fostering dependency. Non-directives, such as Socratic questioning, encourage independence and self-reliance. The N/A category covers utterances unrelated to autonomy.
Self-disclosure in therapy involves a therapist sharing personal experiences or feelings to build trust Marais and McBeath (2021), validate emotions, or model behaviors, enabling a more equitable power relationship in therapy and normalizing clients’ experiences and distress Jolley (2019). It can be categorized as immediate, involving real-time feelings, or non-immediate, sharing past experiences and personal anecdotes Farber (2006). Our framework also includes a None category for utterances lacking self-disclosure.
Metaphor is a symbolic approach that implies similarity between experiences, thoughts, emotions, actions, or objects Wagener (2017). Given their crucial role in fostering learning and comprehension, counselors must be highly skilled at using them effectively in therapy Evans (2010). While some studies show that LLMs struggle with metaphor and analogy comprehension Czinczoll et al. (2022); Tong et al. (2024), we incorporated key concepts from the Conceptual Metaphor Theory Tian et al. (2024), which has shown to enhance LLMs performance in metaphor understanding. We designed our prompts to identify metaphors, along with their corresponding source and target domains. We prompted GPT-4o to analyze therapeutic dynamics in each therapist utterance (see Appendix D.2).
3.4 Linguistic Patterns
We used the Linguistic Inquiry and Word Count (LIWC) tool Boyd et al. (2022) to measure the frequency of words across various linguistic categories within the utterances.